### v3

# Hyperradiance, photon blockade and concurrence in a pair of qubits inside a driven cavity

Abstract
We theoretically study the radiance properties of a pair of qubits inside a single-mode cavity driven by a two-photon drive. Our results show that, when the two qubits are strongly coupled to the cavity field, the collective radiation emitted from the qubits exhibits hyperradiance which can be detected as a signature of two qubit entanglement in the weak-driving regime. We quantify the entanglement in terms of concurrence. Additionally, we study the radiance behaviour in the presence of an intracavity Kerr-nonlinear medium that leads to two-photon blockade. Our results suggest that this system with nonlinearity may act as a quadrature-squeezed and hyperradiant two-photon source.
I Introduction
Cooperative effects of qubits interacting with a single-mode quantized field lead to the well-known phenomena of Dicke superradiance [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and subradiance [15, 15, 7, 5, 16, 17, 18]. Superradiance has been observed in collective emission from two-levels atoms [19, 20, 21], nuclei [22, 23] and pair of ions [24, 7]. In recent times, another cooperative effect called hyperradiance (HR) [25, 26, 27, 28, 29] wherein the emission rate surpasss the superradiant rate of emission has come to the fore. Apart from atomic or nuclear or any massive qubits, the massless photons in the context of cavity quantum electrodynamics (CQED) can also exhibit cooperative effects owing to some nonlinear processes. One such photonic cooperative effect is the photon blockade (PB) [30, 31, 31, 32, 33, 34, 35, 36, 37, 38] which forbids appearance of $(n+1)th$ photon in an $n$-photon blockade. It originates either from intrinsic nonlinearity associated with the dressed-state ladder in the strong-coupling CQED [30, 31, 31, 33, 34, 39] or from the interaction of cavity photons with an intracavity nonlinear medium [32, 37, 38, 36, 40, 41, 42]. Since a cavity field can mediate cooperative effects between a pair of atomic or ionic qubits inside a driven cavity, the question naturally arises if there is any interplay between the radiance and qubit-qubit entanglement or concurrence of the qubit-pair. Furthermore, if the cavity contains a nonlinear medium that can induce multi-photon blockade, then is there any connection of the photon blockade with the radiance and qubit-qubit entanglement? Concurrence is the most reliable measure of two-qubit entanglement, and has been used in the context of quantum cryptography [43], quantum teleportation [44] and generation of quantum random number [45]. But, concurrence has not yet been employed to explicate radiance, to the best of our knowledge.
In this paper, we theoretically propose a two-qubit system coupled to single-mode cavity that is driven by a two-photon drive. A two-photon drive can be realized by pumping an intracavity crystal having second-order nonlinearity with an external laser of frequency twice the cavity frequency [46]. This leads to an optical parametric amplification process and squeezing, whereby a single laser photon generates two cavity photons, thereby breaking the U(1) symmetry of the system. We investigate the radiance of the system using master equation as well as perturbative complex hamiltonian approaches. The results of both approaches are in good agreement in the weak driving regime. We calculate concurrence as a measure of qubit-qubit entanglement.
We quantify the radiance in terms of radiance witness $R$ defined as in Ref. [28]. The negativity of $R$ indicates subradiant regime implying that the radiation in the presence of two qubits is suppressed in comparison to that in the the presence of one qubit. The regime with $R=1$ is termed as superradiant regime. The collective radiation from the qubits characterised by $R>1$ belongs to hyperradiant regime. We demonstrate that, in our model $R$ can reach a value as high as 50 which is substantially higher than the previously reported value [25, 26, 27, 28, 29]. In order to reveal the squeezing property of the radiation we calculate the Wigner distribution [47] of the field. The fact that the Wigner distribution is found to be eliptical in the hyperradiant regime is a clear signature of the quadrature squeezing of the field. This quadrature squeezing is again confirmed by the klyshko’s criterion [48] and the even-odd oscillations of the photon number distributions. We also study the two-photon and three-photon Hanbury Brown-Twiss(HBT) type correlation functions. We find that both the correlations are greater than unity implying super-Poissonian photon statistics.
We next consider an intra-cavity Kerr nonlinear medium-induced two-photon blockade in the system. In our model there exists no PB in the absence of the Kerr medium. The Kerr-nonlinear resonator with a two-photon drive has been studied in a variety of contexts [49, 50, 51, 52]. The conventional PB occurs with a single atomic qubit or a pair of qubits inside a cavity in the strong coupling CQED regime [30, 31]. In the parameter regime of hyperradiance and quadrature squeezing, the intracavity Kerr nonlinear medium leads to the enhancement in the two-photon blockade with finite concurrence. Thus our proposed system with nonlinearity may act as a quadrature-squeezed and hyperradiant two-photon source with concomitant two-qubit entanglement. This suggests that the radiance from a driven cavity containing two qubits may be analyzed for nondestructive detection of two-qubit entanglement.
The rest of this paper is organized as follows: In Sec. II we present our model and methods. In Sec.III we present and discuss our results. Finally, we conclude in Sec. IV.
II The Model and the formulation of the problem
A Hamiltonian
We consider a single-mode cavity with frequency $\omega_{c}$ interacting with a pair of identical qubits with transition frequency $\omega_{a}$. The cavity is driven by a two-photon field
$H_{d}=\eta\left({\hat{a}^{\dagger 2}}e^{-2i\omega_{d}t}+{\hat{a}}^{2}e^{2i\omega_{d}t}\right)$ (1)
where where $\hat{a}(\hat{a^{\dagger}})$ is the annihilation (creation) operator of the cavity field and $\eta$ is the driving strength. The cavity may contain a Kerr medium with nonlinearity strength $\chi$ as shown in Fig. 1. In the frame rotating with the driving frequency $2\omega_{d}$, the Hamiltonian of the system takes the form
$\begin{split}H=\Delta_{c}{\hat{a}}^{\dagger}\hat{a}+\Delta_{a}\sum_{i=1,2}{\hat{\sigma_{i}}}^{\dagger}{\hat{\sigma_{i}}}+(g_{i}{\hat{a}}^{\dagger}{\hat{\sigma_{i}}}+\rm{h.c})+\chi{\hat{a}}^{\dagger}{\hat{a}}^{\dagger}\hat{a}\hat{a}+\eta({\hat{a}^{\dagger 2}}+\hat{a}^{2})\end{split}$ (2)
where $g_{i}=g\cos(2\pi z_{i}/\lambda_{c})$ with $z_{i}$ being the position of the $i$-th atom and $\lambda_{c}$ is the wavelength of the cavity field. Here $\Delta_{a}=\omega_{a}-\omega_{d}$ and $\Delta_{c}=\omega_{c}-\omega_{d}$. The dynamics of the system is governed by the master equation
$\frac{\partial{\hat{\rho}}}{\partial t}=-\frac{i}{\hbar}\left[H,\hat{\rho}\right]+\frac{\kappa}{2}\mathcal{L}_{a}+\sum_{i=1,2}\frac{\gamma_{i}}{2}\mathcal{L}_{\sigma_{i}}$ (3)
where $\rho$ is the density matrix of the system. $\cal L_{\rm{x}}$ is a Liouville super operator where $x$ stands for any of the two operators $a,\sigma_{j}$ ($j=1,2$). Here the Lindbladian superoperators
$\mathcal{L}_{a}$ $=\left[2a\rho a^{\dagger}-\left\{a^{\dagger}a,\rho\right\}\right]~{},$ (4)
$\mathcal{L}_{\sigma_{j}}$ $=\left[2\sigma_{j}\rho\sigma_{j}^{\dagger}-\left\{\sigma_{j}^{\dagger}\sigma_{j},\rho\right\}\right]~{},$
describe the cavity and atomic damping, respectively with $\kappa$ and $\gamma$ being the cavity and atomic decay rates. We solve this master equation numerically[53] and calculate steady-state properties.
Figure 1: Schematic diagram of a pair of identical qubits inside a cavity that may contain a Kerr nonlinear medium. Here, we have shown a pair of two-level atom as qubits. The cavity is driven by a two-photon drive with strength $\eta$ and frequency $\omega_{d}$. Atoms have the transition frequency $\omega_{a}$. The $i$-th atom is coupled to the cavity with strength $g_{i}$ and $\gamma_{i}$ is its decay rate. The decay rate of the cavity is $\kappa$.
B Radiance and Concurrence
One can measure the radiant nature of the two-qubit system based on the correlations between the two qubits using the quantity called radiance witness $R$ [28, 25] defined by
$R=\frac{\langle a^{\dagger}a\rangle_{2}-2\langle a^{\dagger}a\rangle_{1}}{2\langle a^{\dagger}a\rangle_{1}}$ (5)
This involves the intracavity average photon number subject to the existence of one or two qubits inside the cavity. $\langle a^{\dagger}a\rangle_{2}$ denotes the average photon number when both the qubits are present in the cavity. The average photon number when only $i$th single qubit is present in the cavity is denoted by $\langle a^{\dagger}a\rangle_{1}$. $R=0$ implies that the radiation is emitted from uncorrelated qubits, whereas any finite value of $R$ indicates the existance of a correlation between the qubits. The negativity of $R$ implies that the radiation is suppressed compared to that due to two qubits, corresponding to the so-called subradiance regime, while $R>0$ reveals that the radiation is enhanced due to the presence of both qubits. The value of $R=1$ specifically implies that the radiation exhibits scaling proportional to the square of the number of qubits as a characteristic feature of superradiance. When qubits are confined within a cavity, they experience a backaction from the cavity field. This modifies their collective radiative behavior, opening up an intriguing new possibility of hyperradiance regime characterised by $R>1$. We numerically study the properties of radiance solving the master Eq. (3). Furthermore, to gain some analytical insight about hyperradiance, we address the problem by wave function approach with a complex Hamiltonian in the weak driving regime as discussed in Appendix A. In this approach, the mean photon numbers for single- and two-qubit-cavity system in steady-state are given by
$\langle{\hat{a}}^{\dagger}a\rangle_{1}$ $=$ $|C_{g,2}|^{2}+2|C_{e,1}|^{2}$
$\langle{\hat{a}}^{\dagger}a\rangle_{2}$ $=$ $|C_{gg,1}|^{2}+2|C_{gg,2}|^{2},$ (6)
respectively, where $C_{gg,n},C_{g,2},C_{e,1}$ are the probability amplitudes of the states $|gg,n\rangle,|g,2\rangle,|e,1\rangle$ as defined in Appendix A. We thus obtain a perturbative expression of $R$ given by
$\begin{split}R&=\left(\frac{-8\eta^{2}(4g^{2}+16\Delta^{2}+(\kappa+\gamma)^{2})}{|A+iB|^{2}}+\frac{4\eta^{2}(|C|^{2}+8g^{2}|(2\Delta-i\gamma)|^{2})}{|A^{\prime}+iB^{\prime}|^{2}}\right)\times\\ &\frac{D}{8\eta^{2}(4g^{2}+16\Delta^{2}+(\kappa+\gamma)^{2})}\end{split}$ (7)
where
$A$ $=$ $\kappa\gamma+4g^{2}+\kappa^{2}-8\Delta\chi-8\Delta^{2}$
$B$ $=$ $2(\kappa\chi+\Delta\gamma+3\Delta\kappa+\chi\gamma)$
$A^{\prime}$ $=$ $-16\Delta^{3}-16\chi\Delta^{2}+24\Delta g^{2}+2\Delta\kappa^{2}+8\Delta\kappa\gamma+2\Delta\gamma^{2}+8\chi g^{2}+2\chi\kappa\gamma+2\chi\gamma^{2}$
$B^{\prime}$ $=$ $12\Delta^{2}\kappa+12\Delta^{2}\gamma+4\chi\Delta\kappa+12\chi\Delta\gamma-4g^{2}\kappa-8g^{2}\gamma-\kappa^{2}\gamma-\kappa\gamma^{2}$
$C$ $=$ $(8\Delta^{2}-6i\Delta\gamma-2i\Delta\kappa-4g^{2}-\gamma^{2}-\kappa\gamma)$
$D$ $=$ $(4(\kappa+\gamma)^{2}(\Delta+\chi)^{2}+(4g^{2}+\kappa^{2})^{2})+64\Delta(\Delta^{3}+2\Delta^{2}\chi-\Delta g^{2}+\Delta\chi^{2}-g^{2}\chi)+E$
$E$ $=$ $(16\Delta^{2}\kappa^{2}+8g^{2}\kappa\gamma+2\kappa^{3}\gamma+\kappa^{2}\gamma^{2})$
To explore qubit-qubit entanglement we make use of the concurrence [54] defined by
$C(\rho_{q})=[\rm{max}\{{0,{\lambda_{1}}-{\lambda_{2}}-{\lambda_{3}}-{\lambda_{4}}}\}]$ (9)
where ${\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}}$ are the eigenvalues of the matrix $R=\sqrt{\sqrt{\rho_{q}}\tilde{\rho_{q}}\sqrt{\rho_{q}}}$ in decreasing order (the operator $\tilde{\rho_{q}}=(\sigma_{y}\otimes\sigma_{y})\rho_{q}^{*}(\sigma_{y}\otimes\sigma_{y})$ , $\sigma_{y}$ being the Pauli matrix). Here, $\rho_{q}$ is the reduced density matrix of the two qubits. $C$ ranges from zero for unentangled states to one for maximally entangled states.
C Criteria for $n$ photon blockade ($n$PB)
[…]
$P_{n}\geq\Pi_{n},P_{m>n}<\Pi_{m>n}$ (10)
where $\Pi_{n}$ is the Poisson distribution given by $\Pi_{n}=\langle n\rangle^{n}\exp(-\langle n\rangle)/n!$. $\langle n\rangle$ is the average number of photons given by $\langle n\rangle=\rm{Tr}[a^{\dagger}a\rho]$. According to Eq. (10), the $n$PB effect pertains to an increase of the likelihood of $n$ photons while suppressing the probabilities of the photon numbers greater than $n$. The other criterion is based on $g^{(n)}(0)$, an equal-time nth-order correlation function. The $n$PB effect must meet the following requirements
$g^{(n)}(0)\geq 1,g^{(n+1)}(0)<1$ (11)
meaning that the field has $(n+1)$th order sub-Poissonian photon statistics and the $n$th order poissonian or super-Poissonian photon statistics. In particular, the 2PB effect satisfies the correlation functions $g^{(2)}(0)\geq 1$ and $g^{(3)}(0)<1$. The 1PB effect can be characterized by $g^{(2)}(0)<<1$. We have calculated $g^{(2)}(0)$ using the numerical solution of the master Eq. 3 as well as the perturbative analytical solution as obtained in Appendix- A through complex hamiltonian approach. Following the later procedure the equal-time second-order correlation function can be expressed as
$g^{(2)}(0)=\frac{2|C_{gg,2}|^{2}}{(2|C_{gg,2}|^{2}+|C_{+,1}|^{2})^{2}}$ (12)
where $C_{gg,2}$ and $C_{+,1}$ are the probability amplitude of the states $|gg,2\rangle$ and $|+,1\rangle$ as defined in Appendix A.
D Nonclassicality of the field
In order to seek possible connection between $n$-photon blockade and the nonclassical nature of the cavity field, we calculate three quantities, namely squeezing parameter $S_{theta}$, Klyshko’s number $K_{n}$ and radiance witness $R$.
To describe the squeezing character of the cavity field we calculate the squeezing parameter
$S_{\theta}=(\langle X_{\theta}^{2}\rangle-\langle X_{\theta}\rangle^{2})-\frac{1}{2}$ (13)
where $X_{\theta}=(ae^{-i\theta}+a^{\dagger}e^{i\theta})$ is the linear combination of the hermitian quadrature operators $X=\frac{(a+a^{\dagger})}{\sqrt{2}}$ and $Y=\frac{i(a^{\dagger}-a)}{\sqrt{2}}$. Any negative value of $S_{\theta}$ implies squeezing nature of the field characterized by narrowing of the quadrature distribution compared to that of a coherent state. To verify the nonclassicality of the photon number distribution, we use the Klyshko’s criterion [48]
$K_{n}=\frac{(n+1)P_{n-1}P_{n+1}}{nP_{n}^{2}}$ (14)
where $P_{n}$ stands for the photon number distribution. The state is nonclassical when any $K_{n}$ value is below one. This introduces an additional criterion for determining the nonclassical nature of the field apart from $S_{\theta}$. In squeezed states, photons are predominantly distributed among even-numbered states within the Fock space, leading to an absence of odd-numbered photon states. The photon number distribution for squeezed light exhibits oscillations between even and odd numbers. Consequently, the Klyshko’s criterion incorporating odd-even oscillation somewhat mirrors the field’s squeezing characteristics.
E Two-qubit dressed-state picture
In order to explain radiance-concurrence-photon blockade connection, it may be instructive to work in dressed-state picture. In the absence of coherent drive and decay processes, we develop the dressed-state structure for the system in order to have some insight of the physical processes taking place in the system. Throughout our calculations, we have assumed that qubit-cavity coupling for both the atoms are same $g_{1}=g_{2}=g$. We have defined the collective two-qubit basis states as $|gg\rangle$,|±$|\pm\rangle=\frac{|ge>\pm|eg\rangle}{\sqrt{2}}$ and $|ee\rangle$ to describe the dynamics in the dressed-state picture. In the single-photon space the eigenvalues and eigenvectors (dressed states) are obtained as $|\psi_{0}^{(1)}\rangle=-|+,0\rangle$ with $E_{0}^{(1)}=0$ and $|\psi^{(1)}_{\pm}\rangle=\frac{|gg,1\rangle\pm|+,0\rangle}{\sqrt{2}}$ with $E_{\pm}^{(1)}=\pm\sqrt{2}g$. In the two-photon space, the eigenvectors are $|\psi^{(2)}_{0}\rangle=-\frac{1}{\sqrt{3}}|gg,2\rangle+\frac{\sqrt{2}}{\sqrt{3}}|ee,0\rangle$ and $|\phi^{(2)}_{0}\rangle=-|+,1\rangle$ with corresponding eigenvalue $E_{0}^{(2)}=0$. The another eigenvector is $\psi_{\pm}^{(2)}=\frac{1}{\sqrt{3}}|gg,2\rangle\pm\frac{1}{\sqrt{2}}|+,1\rangle+\frac{1}{\sqrt{6}}|ee,0\rangle$ with corresponding eigenvalue $E_{\pm}^{(2)}=\sqrt{6}g$. Due to the symmetric coupling ($g_{1}=g_{2}=g$) the antisymmetric Dicke state $|-,n\rangle$ is decoupled in the $n$photon space ($n=1,2$). In order to know the allowed transitions between the dressed states in the presence of a two-photon dive, we calculate the transition dipole matrix elements between the dressed states as shown in Fig. 2. In the one-photon space, Kerr nonlinearity ($\chi$) will not affect the dressed states. It will act on the states with photon numbers $n\geq 2$. So, the one-photon dressed states and dressed energies in the presence of Kerr nonlinearity will be same as those in the absence of the nonlinearity. In the two-photon space the hamiltonian in the basis $|gg,2\rangle,|\pm,1\rangle$ and $|ee,0\rangle$ is given by,
$H=\begin{pmatrix}2\Delta+2\chi&2g&0&0\\ 2g&2\Delta&0&\sqrt{2}g\\ 0&0&2\Delta&0\\ 0&\sqrt{2}g&0&2\Delta\end{pmatrix}$
. The dressed energies and dressed states are obtained by diagonalizing this Hamiltonian. We have plotted the dressed energies of the system for $\Delta=0,\chi=6\gamma,10\gamma$ in Fig. 3. The dressed energies for $\Delta=0$ with $\chi=6\gamma,g=10\gamma$ are $E_{0}^{(2)}=-21.35,E_{1}^{(2)}=0,E_{2}^{(2)}=3.80,E_{3}^{(2)}=29.55$ and the corresponding dressed states are
$|\psi_{0}^{(2)}\rangle$ $=$ $-0.4472|gg,2\rangle+0.7458|+,1\rangle-0.4938|ee,0\rangle$
$|\psi_{1}^{(2)}\rangle$ $=$ $-|+,1\rangle$
$|\psi_{2}^{(2)}\rangle$ $=$ $-0.5351|gg,2\rangle+0.2193|+,1\rangle+0.8158|ee,0\rangle$
$|\psi_{3}^{(2)}\rangle$ $=$ $-0.7167|gg,2\rangle-0.6291|+,1\rangle-0.3010|ee,0\rangle$ (15)
Figure 2: Dressed-state level diagram for the coupled two-qubit-cavity system without Kerr nonlinearity. Three most significant transitions are shown in this figure. All these three transitions have the same matrix element $\sqrt{\frac{2}{3}}\eta$ with $\eta$ being the pump field Rabi frequency.
Figure 3: Dressed-state energies are plotted as a function of $g$ for $\Delta=0$ with $\chi=6\gamma$ (solid line) and $\chi=10\gamma$ (dashed line).
III Results and discussions
We first present the results in the absence of Kerr medium in subsection III(A) and then discuss the effects of Kerr nonlinearity ($\chi$) in subsection III(B). All the frequency quantities are scaled by the qubit decay constant $\gamma$.
A Generation of hyperradiance by a two-photon drive in the absence of Kerr medium
Here we explore radiance, concurrence between two qubits, squeezing and nonclassicality of the field. Our results suggest a plausible connection between the radiance which is the property of the field and the concurrence that is a measure of entanglement between the qubits. However, there exists no photon blockade in the absence of any Kerr nonlinearity.
The radiance witness $R$ is plotted as a function of the scaled detuning $\Delta/\gamma$ and driving strength $\eta/\gamma$ in Fig. 4 (a). Near $\Delta=0,\pm\sqrt{6}g/2\gamma$ the system exhibits hyperradiant behaviour due to the multiphoton excitations to the dressed eigenstates as illustrated in the dressed-state level diagram of Fig. 2. The strengths of the transition $|\psi^{(0)}\rangle\rightarrow|\psi_{0,\pm}^{(2)}\rangle$ are same and it is $\sqrt{\frac{2}{3}}\eta$. So, unlike [25, 31] where only $|\psi^{(0)}\rangle\rightarrow|\psi_{0}^{(2)}\rangle$ transition was allowed, we have three distinct hyperradiance peaks at three different detunings. In fact, we find regimes where $R\approx 50$ which is substantially greater than the previously reported results so far [28, 26, 27, 25, 29]. To examine whether the enhanced hyperradiance has anything to do with two-qubit entanglement, we plot concurrence between the qubits as a function of $\eta/\gamma$ and $\Delta/\gamma$ for $\chi=0$ in Fig.4 (b). It shows that at $\Delta/\gamma=0,\pm\frac{\sqrt{6}g}{2}$ the two qubits are highly entangled. Now, to verify the nonclassicality, we plot the minimum of the squeezing-parameter Min$[S_{\theta}]$ as a function of $\Delta/\gamma$ and $\eta/\gamma$ in Fig. 4(c). There are three significant peaks at $\Delta=0,\pm\sqrt{6}g/2\gamma$ for weak pumping. The corresponding squeezing angle $\theta_{s}$ is plotted in Fig. 4 (d). The rest of the parameters in Fig. 4 are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$ chosen to fulfill the strong coupling CQED limit. Figures. 5(a) and 5(b) show $R$ and $\langle a^{\dagger}a\rangle$ as a function of $\Delta/\gamma$ at $\eta=0.4$ using both perturbative analytical result given by Eq. (7) and numerically solving master Eq. (3). As shown in these figures, these two results match quite well except slight mismatch in the amplitude or intensity. This is because the states included in the complex hamiltonian approach excludes the gain pathways unlike the master equation approach of Eq. 3. Figure 4(b) can be understood from Fig. 5(c) where two-qubit concurrence $C(\rho_{q})$ is plotted as a function of $\Delta/\gamma$ for $\eta=0.1\gamma$. Three peaks appearing at the transition frequencies as shown in Fig. 5(a) correspond to moderate concurrence. The concurrence goes down when the system is not in the hyperradiant regime or in subradiant regime. Therefore, it is worth noting that the radiance and concurrence seems to have a one-to-one connection as far as the low driving regime is concerned.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: (a) The radiance witness $R$, (b) concurrence $C(\rho_{q})$ and (c) Min[$S_{\theta}$] are plotted as a function of dimensionless pump strength $\eta/\gamma$ and dimensionless detuning $\Delta/\gamma$. (d) Squeezing angle $\theta_{s}$ is plotted on the plane of the $\Delta/\gamma$ and $\eta/\gamma$. The rest of the parameters are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$ and $\chi=0$. The blank areas in (b) corresponds to zero concurrence.
Refer to captionRefer to captionRefer to caption
Figure 5: (a) $R$ and (b) ${{\langle a^{\dagger}a\rangle}_{i}}^{M(C)}$ ($i=1,2$) are plotted as a function of $\Delta/\gamma$ by solving both Eq. (7) (solid) and the master equation (3) (dashed). Here $M(C)$ stands for master equation (complex hamiltonian) and $i$ refers to the number of qubits present in the cavity. (c) Concurrence $C(\rho_{q})$ is plotted as a function of $\Delta/\gamma$. For (a), (b) and (c) $\eta/\gamma=0.1$ and the rest of the parameters are same as in Fig. 4.
We next display the Wigner functions of the cavity field in Figs. 6(a-c) for three typical parameter sets chosen from the Fig. 4 (c) to have strong squeezing. In contrast to the wider range of fluctuations as Fig.6(b) shows, the Wigner functions of Figs. 6 (a) and Fig. 6(c) show elliptical profiles in the parameter regime of squeezing, resulting in the elongation of the distribution in one particular direction. Figures. 6 (a) and 6(c) exhibit squeezing along almost vertical and horizontal directions, respectively, depending on the various pump field detunings.
To illustrate the nonclassical photon number distribution in Fock space, we plot the corresponding Klyshko’s criteria distribution $K_{n}$ in Figs. 6(d–f). The odd-even oscillations of $K_{n}$ alternates up and down mirroring the expected even-odd photon distribution of squeezed states. It is important to note that unlike 6 (b) that corresponds to the resonant case i.e $\Delta=\pm\sqrt{6}g/2\gamma$, the off-resonant cases in Fig. 6(a) and Fig.6(c) show a rapid decay of the even-odd oscillations with the increase of photon number. It means that the squeezed states generated in the off-resonant cases involve mainly two-photon excitations where in resonant cases higher order even number transition processes may take place.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: The density plots of Wigner functions ((a), (b) and (c)) and the logarithm of corresponding Klyshko’s parameter ($\rm{log}K_{n}$) ((d), (e) and (f)) are depicted for three sets of $\eta/\gamma$ and $\Delta/\gamma$ values as mentioned at the top of the each sub-figures of (a), (b) and (c) chosen from Fig. 4(c). The other parameters are same as in Fig. 4
In order to understand the $n$-photon blockade, $\log_{10}g^{(n)}(0)$ ($n=2,3$) is plotted in the plane of $\Delta/\gamma$ and $\eta/\gamma$ in Figs.7 (a) - (b) which show that none of the correlations goes below unity ($\log_{10}g^{(n)}(0)<0$). So it is impossible to find a single-photon or two-photon blockade in this entire parameter range. The reason for the absence of single photon blockade can be explained from the dressed-state picture. The transition from $|\psi^{(0)}\rangle\rightarrow|\psi_{\pm}^{(1)}\rangle$ is not allowed as they are not coupled by two-photon transitions. The value of two and three photon correlations are very large near $\eta=0$. However, in these positions, the average photon numbers are very low making the statistical measurements a difficult task.
Refer to captionRefer to caption
Figure 7: The second and third order HBT correlation functions $\rm{log}_{10}g^{(2)}(0)$ (a) and $\rm{log}_{10}g^{(3)}(0)$ (b) are plotted as a function of $\Delta/\gamma$ and $\eta/\gamma$. The other parameters are same as in Fig. 4. As neither of $g^{(2)}(0)$ and $g^{(3)}(0)$ is less than 1, one- and two- photon blockades do not exist.
So far, we have studied the squeezing and hyperradiance caused by two-photon and multiphoton processes. But, no single or two photon blockade has been found. We next introduce a Kerr nonlinear medium in the cavity to see if we can achieve blockades and how the squeezing and hyperradiance are modified due to the nonlinearity.
B The effect of inra-cavity Kerr-nonlinearity: Photon blockade
Now we introduce Kerr-nonlinearity in order to explore nonlinearity induced photon blockade [32, 55]. In Fig.8 (a) $R$ is plotted as a function of $\eta/\gamma$ and $\Delta/\gamma$. The three peaks obtained correspond to the three transitions in the dressed-state levels. In Fig.8 (b) the significant cocurrence is obtained at $\frac{E_{0}^{(2)}}{2}$, $\frac{E_{2}^{(2)}}{2}$ and $\frac{E_{3}^{(2)}}{2}$ , the same positions where hyperradiance is obtained. However, the introduction of nonlinearity increases the two-qubit concurrence mainly around the resonant transition at $\Delta/\gamma=\frac{E_{2}^{(2)}}{2}$. In Fig.8 (c) Min$[S_{\theta}]$ and (d) $\theta_{s}$ are plotted in the plane of $\eta/\gamma$ and $\Delta/\gamma$ with $\chi=6$. Unlike Fig. 4, this figure shows three peaks at different detunings because the presence of nonlinear medium has modified the dressed state structure. The peaks appear in Fig.8 (a), (b), (c) and (d) are at $\Delta/\gamma=\frac{E_{0}^{(2)}}{2},\frac{E_{2}^{(2)}}{2},\frac{E_{3}^{(2)}}{2}$, respectively. It is because the transitions $|\psi_{0}^{(0)}\rangle\rightarrow\ |\psi_{0}^{(2)}\rangle,|\psi_{2}^{(2)}\rangle,|\psi_{3}^{(2)}\rangle$ are allowed as can be confirmed by calculating the transition dipole matrix elements between the dressed states. It is to be noted that the Min[$S_{\theta}$] value has almost the same minimum as it is obtained for $\chi=0$. The minimum squeezing angle $\theta_{s}$ is usually dependent on the detunings and not on the driving strength as before. To confirm the hyperradiance behaviour in the presence of a Kerr medium, we have plotted $R$ in Fig. 9(a) as well as photon populations in 9(b) as a function of $\Delta/\gamma$. Both perturbative and master equation based results are in good agreement except slight mismatch in the amplitudes as discussed in sec. III(A). We notice that the incorporation of nonlinearity $\chi$ has enhanced the radiance and the population. Next, We have plotted the concurrence $C(\rho_{q}$) as a function of $\Delta/\gamma$ for $\eta=0.1$ in Fig. 9(c). This figure justifies the fact that for detunings that correspond to the resonant transitions between the dressed states, the entanglement is quite significant.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 8: (a) $R$, (b) $C(\rho_{q})$ and (c) Min[$S_{\theta}$] are plotted as a function of $\eta/\gamma$ and $\Delta/\gamma$. The greener area corresponds to relatively stronger squeezing. (d) Squeezing angle $\theta_{s}$ is plotted on the plane of the $\Delta/\gamma$ and $\eta/\gamma$. The white area of figure (b) shows $C(\rho_{q})=0$. The blank areas in both the figure (c) and (d) indicates that the minimum squeezing parameter Min[$S_{\theta}>0$. The other parameters are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$, $\chi=6\gamma$.
Refer to captionRefer to captionRefer to caption
Figure 9: (a) $R$ and (b) ${{\langle a^{\dagger}a\rangle}_{i}}^{M(C)}$ $(i=1,2)$ are plotted as a function of $\Delta/\gamma$ by solving both Eq. (7) (solid) and the master equation (3) (dashed). Here $M(C)$ denotes master equation (complex hamiltonian) and $i$ refers to the number of qubits present in the cavity. (c) Concurrence ($C(\rho_{q})$) is plotted as a function of $\Delta/\gamma$. For (a), (b) and (c) $\eta/\gamma=0.1$ and the rest of the parameters are same as in fig. 8.
Now, to explore the $n$-photon blockade, we plot a phase space curve in Fig. 10 (b) and Fig. 10 (c). In Fig. 10 (b), the yellow region describes the condition when both $g^{(2)}(0)$ and $g^{(3)}(0)$ are less than unity. In this region, we also observed that $g^{(3)}(0)<g^{(2)}(0)$. In Fig. 10 (c) the three yellow regions depict the possibility of two-photon blockade characterised by $g^{(3)}(0)<1,g^{(2)}(0)>1$. In the blue region of (b) and (c), the condition concerning the yellow region is not satisfied. Also, we have ensured that in the yellow region, $\bar{n}$ is appreciable as shown in Fig. 10(a), otherwise the detection of correlation will be a difficult task.
Refer to captionRefer to captionRefer to caption
Figure 10: (a) $\bar{n}$ , (b) the single-photon and (c) two-photon blockade regime (yellow portion) are shown on the plane $\Delta/\gamma$ and $\eta/\gamma$ for the same coupling with $\chi=6\gamma$. The other parameters are same as in Fig. 8.
To ensure the existence of $n$-photon blockade, we choose the most suitable parameters from 10 (b) and (c) and explore the variation as a function of $\Delta/\gamma$. In Fig. 11 (a) and (b) $\bar{n}$ (in the left y axis) and $\log_{10}g^{(n)}(0)$ (in the right y axis) vs. $\Delta/\gamma$ are plotted. In Fig. 11 (a) a strong two-photon blockade regime at $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$ and for $\eta=0.63\gamma$ characterised by $g^{(2)}(0)>1$ and $g^{(3)}(0)<1$ is found. In this regime $g^{(2)}(0)\approx 1$ and $g^{(3)}(0)\approx 0.01$ which indicates an enhanced two-photon blockade. A moderately strong two photon blockade is also found at $\Delta/\gamma=\frac{E_{2}^{(2)}}{2}(g^{(3)}(0)\approx 0.43)$ and $\frac{E_{3}^{(2)}}{2}(g^{(3)}(0)\approx 0.1)$. In Fig. 11 (b) we show single photon blockade characterised by $g^{(3)}(0)<g^{(2)}(0)<1$, observed at $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$ and for $\eta=1.53\gamma$. For the single-photon blockade, $g^{(2)}(0)\approx 0.41$ and $g^{(3)}(0)\approx 0.07$.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 11: (a) and (b) $\bar{n}$ (dashed-dotted), $\rm{log}_{10}g^{(2)}(0)$ (solid) and $\rm{log}_{10}g^{(3)}(0)$ (dotted) are plotted as a function of $\Delta/\gamma$ for $\eta=0.63\gamma$ (a) and $\eta=1.53\gamma$ (b). The other parameters are same as in Fig. 8. (c) and (d) depict deviations of the photon-number distribution from Poisson distribution when $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$. (a) and (c) show two-photon blockade and (b) and (d) show a single-photon blockade at detuning $\frac{E_{0}^{(2)}}{2}$.
The foregoing results can also be confirmed by comparing the photon-number distributions with the Poisson distribution. In order to verify the two-photon blockade, we plot $\frac{P_{n}-{\Pi}_{n}}{\Pi_{n}}$ with respect to photon number in Fig. 11 (c) which shows that only two-photon probabilities are enhanced ($P_{2}>\Pi_{2}$) and all the higher order photon probabilities are suppressed ($P_{n}<\Pi_{n}$ for $n>2$). To verify single-photon blockade in (b) we plot the same as (c) in Fig. 11 (d) and notice that only single-photon probabilities are enhanced ($P_{1}>\Pi_{1}$) and all the higher order photon probabilities are suppressed ($P_{n}<\Pi_{n}$ for $n>1$). It is to be noted that in both the cases (a) and (b), we plot $\bar{n}$ to confirm the presence of appreciable photon number at the respective points. The rest of the parameters are same as in Fig. 10.
Now, to further corroborate our results, we have judiciously chosen a parameter regime and explored all the statistical properties. In Fig. 12 (a) and (b), we have plotted $R$ and $C(\rho_{q})$ as a function of $\Delta/\gamma$ and $\eta/\gamma$. We show that at $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$ the highest peak in the $R$ vs. $\Delta/\gamma$ curve has a finite concurrence. We also investigate squeezing in the same regime. The Wigner distribution in Fig. 12 (c) shows eliptical profile in the X-Y space in the diagonal direction that indicates the emitted radiation is squeezed. Squeezing is indeed verified by the Kleysko’s parameter $K_{n}$ in Fig. 12 (d) which shows prominent even- odd oscillations in the photon number space. Lastly, we plot $\rm{log}_{10}g^{(n)}(0)$ (right y-axis) and $\bar{n}$ (left y-axis) as a function of $\Delta/\gamma$ in Fig. 12 (e). $\bar{n}$ shows three distinct peaks at $\frac{E_{0}^{(2)}}{2},\frac{E_{2}^{(2)}}{2},\frac{E_{3}^{(2)}}{2}$. In these positions $g^{3}(0)<1$ and $g^{2}(0)>1$ indicating a 2P blockade. In these three positions the $g^{2}(0)\approx 2,23.5,3$ and the $g^{(3)}(0)$ is achieved as low as $g^{(3)}(0)\approx 0.05,0.6,0.02$. So, $g^{(2)}(0)$ is two order of magnitude higher than $g^{(3)}(0)$ at $\frac{E_{0}^{(2)}}{2}$ which indicates an enhanced 2P blockade. The strong 2P blockade is again confirmed by plotting $\frac{P_{n}-\Pi_{n}}{\Pi_{n}}$ as a function of photon number. This quantity shows a peak at $n=2$ ($P_{2}>\Pi_{2}$) and its values for all higher photon numbers are negative ($P_{n}<\Pi_{n}$ for $n>2$) indicating a strong two-photon emission with suppression of higher photon number states. To further corroborate the numerical results we have plotted $\rm{log}_{10}g^{(2)}(0)$ as a function of $\Delta/\gamma$ using both the Eq. (12) and Eq. (3) in Fig. 13 for the same parameter regime as in Fig. 12. Both the results are in good agreement except slight mismatch in the amplitudes for the same reason as mentioned earlier. Thus our system can act as a hyperradiant two-photon squeezed light generator.
Figure 12: (a) $R$ and (b) $C(\rho_{q})$ are plotted as a function of $\Delta/\gamma$. (c) Wigner function is plotted in $X$ and $Y$ space, (d) logarithm of the Kleyshko’s parameter $K_{n}$ is plotted as a function of $n$. (e) $\bar{n}$ (solid), $\rm{log}_{10}g^{(n)}(0)$ ($n=2$ $(\rm{dashed})$, $3$ $(\rm{dashed-dotted})$) vs. $\Delta/\gamma$ and (f) $\frac{P_{n}-\Pi_{n}}{\Pi_{n}}$ vs. $n$ are shown. For all these figures $\eta=0.4\gamma$. For (c), (d) and (f) $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$. The rest of the parameters are same as in Fig. 8.
Figure 13: (a) $\rm{log}_{10}g^{(2)}(0)$ is plotted as a function of $\Delta/\gamma$ for $\eta=0.4\gamma$ by solving both Eq. (12) (solid) and the Master equation Eq.(3) (dashed). The rest of the parameters are same as in Fig. 8.
IV Conclusions
In summary, we have studied the quantum properties of a CQED system with a pair of qubits inside a cavity driven by a two-photon drive. We have shown that the qubits give rise to hyperradiance which can be linked with the two-qubit entanglement in low driving regime. We have further demonstrated that the emitted hyperradiant light can be squeezed in quadrature. To confirm squeezing, we have calculated Wigner distribution of the field and Kleyshko’s parameter. The Wigner distribution shows eliptical profile while the Kleyshko’s parameter exhibits even-odd oscillations as a clear signature of squeezing. We have also explored the influence of an intra-cavity Kerr-nonlinear medium on the system. The nonlinear medium induces photon blockade that can be employed to control the photon statistics of the field. We have shown that the hyperradiant field can be generated with single as well as two-photon blockades in suitable parameter regimes. The hyperradiant field with two-photon blockade is found to be quadrature-squeezed. Interestingly in these regimes, the system possesses a finite two-qubit concurrence confirming the entanglement between the qubits. Our findings may have potential applications in quantum communication and quantum networking.
acknowledgement
AD is thankful to Subhanka Mal for fruitful scientific discussions.
Appendix A Analysis using non-hermitian Hamiltonian approach
In what follows, we present approximate analytical expressions for the radiance $R$ of the transmitted field. Under the scenario of weak driving ($\eta<<g,\kappa$) the total excitation of the qubit-cavity system can be assumed to be two as in Refs. [56, 35]. Under the weak driving limit, the qubit-cavity system remains predominantly in the ground state. So, the terms arising as $\kappa a\rho a^{\dagger}$ and $\gamma\sigma_{j}\rho\sigma_{j}^{\dagger}$ in the Lindblad master equation can safely be ignored [57, 58]. Under these circumstance, the joint wave function of the pair of qubits and the cavity field can be reasonably approximated in the two-photon manifold with the ansatz
$\begin{split}|\psi_{2}(t)\rangle&=C_{gg,0}(t)|gg,0\rangle+C_{gg,2}(t)|gg,2\rangle+C_{gg,1}(t)|gg,1\rangle+C_{+,1}(t)|+,1\rangle+\\ &C_{+,0}(t)|+,0\rangle+C_{ee,0}(t)|ee,0\rangle.\end{split}$ (16)
and the corresponding Hamiltonian will be
$\begin{split}H^{2}_{nh}&=(\Delta_{c}-i\kappa/2)\hat{a}^{\dagger}\hat{a}+(\Delta_{a}-i\gamma_{1}/2)\hat{\sigma_{1}}^{\dagger}\hat{\sigma_{1}}+(\Delta_{a}-i\gamma_{2}/2)\hat{\sigma_{2}}^{\dagger}\hat{\sigma_{2}}+\\ &g_{1}(\hat{a}^{\dagger}\hat{\sigma_{1}}+\hat{a}\hat{\sigma_{1}}^{\dagger})+g_{2}(\hat{a}^{\dagger}\hat{\sigma_{2}}+\hat{a}\hat{\sigma_{2}}^{\dagger})+\chi{\hat{a}}^{\dagger}{\hat{a}}^{\dagger}\hat{a}\hat{a}+\eta({\hat{a}^{\dagger^{2}}}+\hat{a}^{2}).\end{split}$ (17)
The coefficient $C_{lm,n}$ stands for the probability amplitude of the corresponding state $|lm,n\rangle$.
We also treat single qubit-cavity system in the two photon space with the ansatz wave function
$\begin{split}|\psi_{1}(t)\rangle=C_{g,0}(t)|g,0\rangle+C_{g,1}(t)|g,1\rangle+C_{g,2}(t)|g,2\rangle+C_{e,0}(t)|e,0\rangle+C_{e,1}(t)|e,1\rangle\end{split}$ (18)
and the non-Hermitian Hamiltonian is given by
$H^{1}_{nh}=(\Delta_{c}-i\kappa/2)\hat{a}^{\dagger}\hat{a}+(\Delta_{a}-i\gamma/2)\hat{\sigma_{1}}^{\dagger}\hat{\sigma_{1}}+g(\hat{a}^{\dagger}\hat{\sigma_{1}}+\hat{a}\hat{\sigma_{1}}^{\dagger})+\chi{\hat{a}}^{\dagger}{\hat{a}}^{\dagger}\hat{a}\hat{a}+\eta({\hat{a}^{\dagger^{2}}}+\hat{a}^{2}).$ (19)
The coefficient $C_{m,n}$ stands for the probability amplitude of the corresponding state $|m,n\rangle$. Time dependent Schroedinger equation incorporating $H^{1}_{nh}$ and $\psi_{1}(t)$ yields a set of coupled differential equations as follows
$i\hbar\dot{C_{g,1}}$ $=$ $\Delta C_{g,1}+gC_{e,0}-i\frac{\kappa}{2}C^{1}_{g,1}$
$i\hbar\dot{C_{g,2}}$ $=$ $2\Delta C_{g,2}+\sqrt{2}gC_{e,1}+2\chi C_{g,2}+\sqrt{2}\eta C_{g,0}-i{\kappa}C^{1}_{g,2}$
$i\hbar\dot{C_{e,0}}$ $=$ $\Delta C_{e,0}+gC_{g,1}-i{\gamma}C_{e,0}$
$i\hbar\dot{C_{e,1}}$ $=$ $2\Delta C_{e,1}+\sqrt{2}gC{g,2}-i(\frac{\kappa}{2}+\frac{\gamma}{2})C^{1}_{e,1}.$ (20)
In the weak driving limit, we have the relationship $C_{g,0}>>C_{g,2},C_{e,1}>>C_{g,1},C_{e,0}$ since it is a two photon drive. So, we assume that $C_{g,0}\approx 1$. . Again using Eqs. 17 and 16, we get the coupled differential equations of the form
$i\hbar\dot{C_{gg,2}}$ $=$ $2\Delta C_{gg,2}+2gC_{+,1}+2\chi C_{gg,2}+\sqrt{2}\eta C_{gg,0}-i\kappa C_{gg,2}$
$i\hbar\dot{C_{gg,1}}$ $=$ $\Delta C_{gg,1}+\sqrt{2}gC_{-,0}+\sqrt{6}\eta C_{gg,3}-\frac{i\kappa}{2}C_{gg,1}$
$i\hbar\dot{C_{+,1}}$ $=$ $2\Delta C_{+,1}+\sqrt{2}gC_{ee,0}+2gC_{gg,2}+2\chi C_{+,2}+\sqrt{2}\eta C_{+,0}-i(\frac{\kappa}{2}+\frac{\gamma}{2})C_{+,1}$
$i\hbar\dot{C_{+,0}}$ $=$ $\Delta C_{+,0}+\sqrt{2}gC_{gg,1}-i\frac{\gamma}{2}C_{+,0}$
$i\hbar\dot{C_{ee,0}}$ $=$ $2\Delta C_{ee,0}+\sqrt{2}gC_{+,1}-i\gamma C_{ee,0}$ (21)
Again, considering weak driving we assume that $C_{gg,0}\approx 1$. We solve these dynamical equations in the steady-state. The solution of Eq.(20) and Eq.(21) are given by,
$C_{g,2}$ $=$ $\frac{-\sqrt{2}\eta}{2((\Delta_{k}+\chi)(\Delta_{k}+\Delta_{l}))-g^{2}}$
$C_{e,1}$ $=$ $\frac{\eta g}{2((\Delta_{k}+\chi)(\Delta_{k}+\Delta_{l}))-g^{2}}$ (22)
$C_{gg,2}$ $=$ $\frac{-\sqrt{2}\eta(\Delta_{l}^{2}+\Delta_{k}\Delta_{l}-g^{2})}{2(\Delta_{l}(\Delta_{k}+\chi)(\Delta_{k}+\Delta_{l}))-g^{2}(\Delta_{k}+2\Delta_{l}+\chi)}$
$C_{+,1}$ $=$ $\frac{-\sqrt{2}\Delta_{l}\eta g}{2(\Delta_{l}(\Delta_{k}+\chi)(\Delta_{k}+\Delta_{l}))-g^{2}(\Delta_{k}+2\Delta_{l}+\chi)}$ (23)
For the sake of simplicity, we have defined two effective complex detunings as $\Delta_{\kappa}=\Delta-\frac{i\kappa}{2}$ and $\Delta_{\gamma}=\Delta-\frac{i\gamma}{2}$.